OpenStudy (anonymous):

State the horizontal asymptote of the rational function. f(x) = (9x^2 - 3x - 8)/(4x^2 - 5x + 3)

OpenStudy (anonymous):

no like ive tried the problem a dozen times at least i just cant get the answer to be one of the choices...:(

OpenStudy (anonymous):

y = 3/5 y = 9/4 y = 0 None

OpenStudy (anonymous):

those are the choices

OpenStudy (luigi0210):

Well did you read the page? It explains it pretty well.

OpenStudy (luigi0210):

In this example n=m, so that means the Horiz Asym: a/b

OpenStudy (anonymous):

yes i read it but im not the math brain so i need help

OpenStudy (luigi0210):

When dealing with asym. you have to look at the exponents. So tell me are the exponents the same power?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

Good, so in case instead of it being that long equation it turns into this: \[\frac{ 9x^2-3x-8 }{ 4x^2-5x+3 }=\frac{9x^2}{4x^2}=\frac{9}{4}\]

OpenStudy (anonymous):

oooooooh!!! i get it so its only the one with the exponent x's?

OpenStudy (luigi0210):

If they have the same power. And only if those powers are also the higher ones.

OpenStudy (anonymous):

ok i get it! so it wouldnt work if it was like 9x^3 and 4x^2? and it would if both were x^5?

OpenStudy (anonymous):

(just examples to make sure)

OpenStudy (luigi0210):

Here I'll show you: \[\frac{x^2}{x^2}=\frac{1}{1}\] \[\frac{x^3}{x^2}=slant \] \[\frac{x}{x^2}= 0 \]

OpenStudy (anonymous):

ok thanks! :)

OpenStudy (luigi0210):

No problem.